OpenStudy (anonymous):

The diagonal of a square A is (x+y). The diagonal of a square B with twice the area of A is

OpenStudy (anonymous):

if you had a square that was 4 by 4 that are would be 16 the diagonal would be 4^2 + 4^2 = c^2 16 + 16 = c^2 c = sqrt(32) c = 4sqrt(2) a square with double that area would have an area of 32 each side would be sqrt(32) = 4sqrt(2) the diagonal would be (4sqrt(2))^2 + (4sqrt(2))^2 = c^2 32 + 32 = c^2 c = sqrt(64) c = 8 the ratio is 8 -- 4sqrt(2) which is sqrt(2) larger so sqrt(2)(x + y)

OpenStudy (anonymous):

correct answer